# DetectRL: Benchmarking LLM-Generated Text Detection in Real-World Scenarios

### v1

@@ Abstract
Detecting text generated by large language models (LLMs) is of great recent interest. With zero-shot methods like DetectGPT, detection capabilities have reached impressive levels. However, the reliability of existing detectors in real-world applications remains underexplored. In this study, we present a new benchmark, DetectRL, highlighting that even state-of-the-art (SOTA) detection techniques still underperformed in this task. We collected human-written datasets from domains where LLMs are particularly prone to misuse. Using popular LLMs, we generated data that better aligns with real-world applications. Unlike previous studies, we employed heuristic rules to create adversarial LLM-generated text, simulating advanced prompt usages, human revisions like word substitutions, and writing errors. Our development of DetectRL reveals the strengths and limitations of current SOTA detectors. More importantly, we analyzed the potential impact of writing styles, model types, attack methods, the text lengths, and real-world human writing factors on different types of detectors. We believe DetectRL could serve as an effective benchmark for assessing detectors in real-world scenarios, evolving with advanced attack methods, thus providing more stressful evaluation to drive the development of more efficient detectors [^1] Data and code are publicly available at: https://github.com/NLP2CT/DetectRL.
@@ 1 Introduction
Detecting text generated by LLMs is a challenging task. It is often more difficult for humans than for detection techniques to identify LLM-generated text, as humans typically underperform detection methods designed for this purpose [1]. Recently, the implications of LLM-generated content have come into focus, highlighting their significant societal and academic impacts and associated risks [2, 3]. The main concerns stem from the hallucinations and misuse of LLMs [4], leading to issues such as plagiarism [5], the spread of fake news [6], and challenges to educators and human scholarship in AI-assisted academic writing [7]. Previous and current popular detection benchmarks, such as TuringBench [8], MGTBench [9], MULTITuDE [10], MAGE [11] and M4 [12], have primarily focused on evaluating detectors’ performance across various domains, generative models, and languages by constructing idealized test data. However, they have overlooked the assessment of detectors’ capabilities in more common scenarios encountered in practical applications [4], such as various prompt usages and human revisions, as shown in Table 1.
@@
Domains LLMs Prompts Revision Errors Mixing Generalization Length Length Human Writing
@@
In this paper, we study the following questions: (1) How do SOTA LLM-generated text detectors perform in real-world application scenarios? (2) What real-world factors influence detector performance, and to what extent? We investigate these questions by introducing DetectRL, a novel benchmark for LLM-generated text detection. We achieve this by crafting challenges that are commonly encountered in real-world scenarios. These challenges simulate various prompts usage of human, human revisions of text such as word substitutions, and common writing errors, including spelling mistakes. To enhance these simulations, we incorporate well-designed attack methods like prompt-based attacks, paraphrasing, adversarial perturbations, and data mixing. We selected data from domains where LLMs are frequently used and prone to abuse, such as academic writing, news writing, creative writing, and social media, to serve as samples of human-written text. To create LLM-generated texts that closely resemble real-world application scenarios, we employed powerful and widely used LLMs, including GPT-3.5-turbo [13], PaLM-2-bison [14], Claude-instant [15], and Llama-2-70b [16]. Furthermore, to ensure a wider diversity of text length, we filtered out shorter texts and applied a varying length augmentation method. This approach significantly broadened the range of text lengths available for detection, enhancing the practical value of the task. We balanced sample distributions across domains, LLMs, and attack types in all test scenarios to enhance diversity, thereby creating more challenging evaluations. These distribution variances are common in real-world scenarios but are often overlooked in ideal test environments where current detectors are developed.
The construction of this benchmark was highly effective in achieving our goals. The experimental results present a significant challenge to existing detection methods. Current detectors, particularly those employing zero-shot techniques, often struggle with accurately identifying LLM-generated texts. For example, adversarial perturbation attacks reduce the performance of all zero-shot detectors by an average of 34.48% AUROC. In contrast, supervised detectors have demonstrated robust detection capabilities in various domains, generative models, and attacks settings.
@@ 2.1 Framework
Perturbation attacks mainly involve introducing adversarial perturbations on text directly generated by LLMs. These attacks can effectively simulate common writing errors, word substitutions or other noises in real-world applications. We utilize TextFooler [24] for Word-level Perturbations, DeepWordBug [25] for Character-level Perturbations, and TextBugger [25] for Sentence-level Perturbations, all implemented using TextAttack [26].
@@ 2.2 Task definition
This task aims to evaluate the foundational performance of detectors in different domains, generators, and attack strategies, focusing specifically on their in-domain performance in various real-world scenarios. We use the average performance score as the metric for foundational in-domain capabilities assessment.
@@
This task evaluates how real-world human writing factors impact the performance of detectors. In this innovative assessment, we simulate and replicate these factors like spelling errors by applying attacks on human-written texts, highlighting the challenges they pose to detectors in real-world scenarios.
@@ 2.3 Benchmark statistics
The statistics for the collected data are presented in Appendix Table 9. This dataset includes a total of 100,800 human-written samples, with 11,200 being raw samples and 89,600 modified through attack manipulations. Additionally, there are 134,400 samples generated by LLMs, categorized as follows: 11,200 with direct prompt , 22,400 with prompt attacks, 33,600 with paraphrase attacks, 33,600 with perturbation attacks, and 22,400 with data mixing. For the assessment task we designed, we extracted relevant data from this dataset to develop the DetectRL benchmark. Detailed statistics for each task in the benchmark are presented in Table 2. Samples were carefully selected from the constructed dataset to align with task design, ensuring a balance across domains, LLMs, and attacks. The training data was specifically tailored for both supervised and zero-shot detectors, and performance was evaluated using common test sets. The statistics of the textual features of DetectRL samples are illustrated in Figure 2. For a more detailed analysis, please refer to Appendix subsection D.6.
@@ 2.4 Evaluation metrics
We employ AUROC and $F_{1}\text{ Score}$ as the main evaluation metrics. AUROC is widely used for assessing zero-shot detection methods [27] because it considers the True Positive Rate (TPR) and False Positive Rate (FPR) across different classification thresholds. This makes AUROC particularly useful for evaluating detector performance at different thresholds. The $F_{1}\text{ Score}$ provides a comprehensive evaluation of detector capabilities by balancing the model’s Precision and Recall. Additionally, we provide detailed Precision and Recall scores in Appendix F for further reference, with a specific focus on Recall to highlight the detectors’ effectiveness in identifying LLM-generated text.
@@ 3 Experiments and discussion
In this section, we organize our experiments and discussions from five distinct perspectives: (1) Benchmarking the cutting-edge detectors: We evaluate the current SOTA detectors against our benchmark to identify ongoing challenges. (2) Robustness analysis: We examine factors contributing to robustness issues in various domains, LLMs, and attack scenarios. (3) Assessing generalization: We investigate how well detectors perform on data they were not specifically trained on, highlighting their out-of-distribution robustness. (4) Length discrimination: We examine the ability of detectors to differentiate between texts of varying lengths and the impact of training on such texts. (5) Real-world human writing scenarios: We analyze the effects of real-world post-processing and mistake in human-written texts, discussing their implications to provide more nuanced and valuable insights.
@@ 3.1 Benchmarking detectors
We employ a variety of SOTA detectors to assess the difficulty of DetectRL. Given that LLMs in real-world scenarios are often black-box and inaccessible, we exclude watermarking methods from our evaluation. Our evaluation encompasses prominent zero-shot techniques and supervised fine-tuned classifiers, including Log-Likelihood [28], Entropy [29], Rank [30], Log-Rank [30], LRR [31], NPR [31], DetectGPT [27], Fast-DetectGPT [32], Revise-Detect. [33], DNA-GPT [34], Binoculars [35], RoBERTa Classifier (RoB [36]), and XLM-RoBERTa Classifier [^4] Please refer to Appendix E.2 for comprehensive detectors specifications. (X-RoB [37]). For the white-box zero-shot detection method, we employ the GPT-Neo-2.7B [38] as the scoring model, in line with the methodology proposed in Fast-DetectGPT [32], to detect the text generated by black-box LLMs. For the black-box zero-shot detection method like Revise-Detect. [33] and DNA-GPT [34], we use GPT-4o-Mini [39] to perform operations such as text revision and text continuation. For supervised detectors, all classifiers are trained using the same parameters. For detailed training parameters, please refer to Table 12 and subsection E.2.
@@
Table 3: The overall leaderboard for LLM-generated text detectors in real-world scenarios ranks detectors based on their robustness and generalization on multiple domains, LLMs, and attack scenarios. Rankings also consider the impact of text length in training-time and test-time phase, as well as performance against real-world human writing factors.
@@
Detectors $\downarrow$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ $F_{1}$ $F_{1}$ $F_{1}$ $F_{1}$ $F_{1}$ AUROC $F_{1}$ $F_{1}$
@@
Log-Likelihood - 65.35 57.55 45.68 41.32 68.00 59.38 75.84 67.22 63.22 56.37
Entropy - 48.39 29.71 67.84 57.23 39.06 20.55 28.82 02.14 46.53 27.66
Rank - 57.17 54.62 36.87 22.47 56.26 50.90 55.08 51.90 51.09 44.97
Log-Rank - 67.01 60.09 46.74 42.60 67.58 57.57 76.40 69.88 64.43 57.78
LRR - 70.54 61.34 50.09 38.38 64.65 53.09 76.61 68.99 65.47 55.70
NPR - 53.85 49.65 34.59 18.31 54.96 52.30 50.09 45.39 48.87 41.16
DetectGPT - 22.15 00.00 12.21 00.00 58.95 50.83 44.43 35.25 34.44 21.02
DNA-GPT - 67.41 58.30 64.22 45.09 69.04 58.25 78.17 69.28 69.71 57.23
Revise-Detect. - 70.40 37.51 50.34 46.07 73.24 64.29 75.01 68.71 67.75 54.65
@@
Log-Likelihood - 62.89 57.80 43.32 28.10 70.03 60.73 75.65 65.90 62.47 53.63
Entropy - 46.84 23.29 52.25 30.42 45.34 16.56 43.48 66.75 46.98 34.26
Rank - 52.19 49.32 41.68 22.78 50.40 41.74 57.05 54.40 50.33 42.56
Log-Rank - 62.84 56.87 43.32 30.12 70.89 63.09 77.97 66.66 63.76 54.68
LRR - 61.61 52.12 43.30 18.91 71.17 65.51 83.65 75.51 64.43 53.01
NPR - 50.29 43.81 41.64 32.91 44.64 34.77 52.53 48.68 47.78 40.54
DetectGPT - 43.46 26.27 32.86 12.56 26.72 00.00 36.71 20.40 34.44 14.81
DNA-GPT - 61.87 55.04 48.88 25.67 71.48 60.77 75.22 62.89 64.86 51.59
Revise-Detect. - 70.10 62.72 49.87 27.28 69.84 59.03 75.65 65.87 66.87 53.73
Binoculars - 88.14 82.50 55.15 39.35 93.30 88.20 96.64 92.30 83.31 75.59
Fast-DetectGPT - 65.56 59.55 30.01 00.00 65.99 57.58 76.79 69.08 59.59 46.55
@@
Significant Challenge
Our benchmarks reveal significant challenges in the current LLM-generated text detection research. We found that incorporating a mix distribution of domains, LLMs, and attack types increases the testing pressure of zero-shot methods. For example, in the multi-LLM setting, the average AUROC of all zero-shot detectors is only 58.61%. This is because data from each LLM spans various domains and attack methods, leading to substantial distribution differences even within data from the same LLM. These variations are often overlooked in ideal testing environments, making it difficult for zero-shot detectors developed based on them to work effectively. Specifically, zero-shot detectors struggle against powerful LLMs, achieving an average AUROC of only 72.80% on texts generated via direct prompting, with no method surpassing a 90% AUROC. The performance of these detectors declines markedly under well-designed attacks that simulate real-world scenarios, with average decreases of 2.59% in prompt attacks, 17.50% in paraphrase attacks, 38.48% in perturbation attacks, and 20.29% in data mixing scenarios. In contrast, supervised methods demonstrate impressive effectiveness, achieving an average AUROC of 99.42% on data generated through direct prompting and maintaining robustness against well-designed attacks.
Unexpectedly, recent advancements in LLM-generated text detection, such as DetectGPT [27], NPR [31], Fast-DetectGPT [32], and DNA-GPT [34], did not perform as expected on our benchmark. Their performance was even weaker than some traditional zero-shot baselines. Analysis across various domains and LLMs revealed a general lack of robustness as a potential underlying issue. For instance, DetectGPT’s performance was notably low, with only 22.15% AUROC in academic writing (ArXiv) and 12.21% AUROC in news writing (Xsum), though it achieved 58.95% AUROC in creative writing and 51.82% AUROC in social media (Yelp Review). A similar trend was observed with the best zero-shot detector, Binoculars, which performed more than 10% lower in academic writing and news writing compared to other domains. Additionally, Binoculars showed significantly reduced effectiveness on text generated by Claude, achieving only 55.15% AUROC, while presenting 88.14%, 93.30%, and 96.64% AUROC on text generated by GPT-3.5, PaLM-2, and Llama-2, respectively. These findings suggest that the performance disparities of detectors across different domains and LLMs are significantly exacerbated when subjected to well-designed attacks, contradicting the conclusions of some studies such as [32].

### v3

@@ Abstract
Detecting text generated by large language models (LLMs) is of great recent interest. With zero-shot methods like DetectGPT, detection capabilities have reached impressive levels. However, the reliability of existing detectors in real-world applications remains underexplored. In this study, we present a new benchmark, DetectRL, highlighting that even state-of-the-art (SOTA) detection techniques still underperformed in this task. We collected human-written datasets from domains where LLMs are particularly prone to misuse. Using popular LLMs, we generated data that better aligns with real-world applications. Unlike previous studies, we employed heuristic rules to create adversarial LLM-generated text, simulating various prompts usages, human revisions like word substitutions, and writing noises like spelling mistakes. Our development of DetectRL reveals the strengths and limitations of current SOTA detectors. More importantly, we analyzed the potential impact of writing styles, model types, attack methods, the text lengths, and real-world human writing factors on different types of detectors. We believe DetectRL could serve as an effective benchmark for assessing detectors in real-world scenarios, evolving with advanced attack methods, thus providing more stressful evaluation to drive the development of more efficient detectors. [^1] Data and code are publicly available at: https://github.com/NLP2CT/DetectRL
@@ 1 Introduction
Detecting text generated by LLMs is a challenging task. It is often more difficult for humans than for detection techniques to identify LLM-generated text, as humans typically underperform detection methods designed for this purpose [1]. Recently, the implications of LLM-generated content have come into focus, highlighting their significant societal and academic impacts and associated risks [2, 3]. The main concerns stem from the hallucinations and misuse of LLMs [4], leading to issues such as plagiarism [5], the spread of fake news [6], and challenges to educators and human scholarship in AI-assisted academic writing [7]. Previous and current popular detection benchmarks, such as TuringBench [8], MGTBench [9], MULTITuDE [10], MAGE [11] and M4 [12], have primarily focused on evaluating detectors’ performance across various domains, generative models, and languages by constructing idealized test data. However, they have overlooked the assessment of detectors’ capabilities in more common scenarios encountered in practical applications [4], such as various prompt usages, human revisions, and writing noises, as shown in Table 1.
@@
Domains LLMs Prompts Revision Noises Mixing Generalization Length Length Human Writing
@@
In this paper, we study the following questions: (1) How do SOTA LLM-generated text detectors perform in real-world application scenarios? (2) What real-world factors influence detector performance, and to what extent? We investigate these questions by introducing DetectRL, a novel benchmark for LLM-generated text detection. We achieve this by crafting challenges that are commonly encountered in real-world scenarios. These challenges simulate various prompts usages of human, human revisions of text such as word substitutions, and adversarial writing noises, including spelling mistakes. To enhance these simulations, we incorporate well-designed attack methods like prompt-based attacks, paraphrasing, adversarial perturbations, and data mixing. We selected data from domains where LLMs are frequently used and prone to abuse, such as academic writing, news writing, creative writing, and social media, to serve as samples of human-written text. To create LLM-generated texts that closely resemble real-world application scenarios, we employed powerful and widely used LLMs, including GPT-3.5-turbo [13], PaLM-2-bison [14], Claude-instant [15], and Llama-2-70b [16]. Furthermore, to ensure a wider diversity of text length, we filtered out shorter texts and applied a varying length augmentation method. This approach significantly broadened the range of text lengths available for detection, enhancing the practical value of the task. We balanced sample distributions across domains, LLMs, and attack types in all test scenarios to enhance diversity, thereby creating more challenging evaluations. These distribution variances are common in real-world scenarios but are often overlooked in ideal test environments where current detectors are developed.
The construction of this benchmark was highly effective in achieving our goals. The experimental results present a significant challenge to existing detection methods. Current detectors, particularly those employing zero-shot techniques, often struggle with accurately identifying LLM-generated texts. For example, adversarial perturbation attacks reduce the performance of all zero-shot detectors by an average of 39.28% AUROC. In contrast, supervised detectors have demonstrated robust detection capabilities in various domains, generative models, and attacks settings.
@@ 2.1 Framework
Perturbation attacks mainly involve introducing adversarial perturbations on text directly generated by LLMs. These attacks can effectively simulate common writing errors, character or word substitutions or other adversarial noises in real-world applications. We utilized DeepWordBug [24] for Character-level Perturbations, TextFooler [25] for Word-level Perturbations, and TextBugger [24] for Sentence-level Perturbations, all implemented using TextAttack [26].
@@ 2.2 Task definition
This task aims to evaluate the foundational performance of detectors in different domains, generators, and attack strategies, focusing specifically on their in-domain robustness in various real-world scenarios. We use the average performance score as the assessment metric.
@@
This task evaluates how real-world human writing factors impact the performance of detectors. In this innovative assessment, we simulate and replicate these factors like word substitutions and spelling errors by applying attacks on human-written texts, highlighting the challenges they pose to detectors in real-world scenarios.
@@ 2.3 Benchmark statistics
The statistics for the collected data are presented in Appendix Table 9. This dataset comprises 100,800 human-written samples, including 11,200 raw samples and 89,600 samples modified via attack manipulations. Additionally, it contains 134,400 samples generated by LLMs, categorized as follows: 11,200 samples generated with direct prompt, 22,400 with prompt attacks, 33,600 with paraphrase attacks, 33,600 with perturbation attacks, and 22,400 with data mixing. To evaluate detectors performance, we designed the DetectRL benchmark by carefully extracting relevant subsets of data to align with the task design. The selected samples ensure a balance across domains, LLMs, and attack types. The training data was specifically tailored for both supervised and zero-shot detectors, and performance was evaluated using common test sets. Detailed statistics for each task and the analysis of the textual features of DetectRL samples are presented in Table 2. For a more detailed analysis, please refer to Appendix D.6.
@@ 2.4 Evaluation metrics
We employ AUROC and $F_{1}\text{ Score}$ as the main evaluation metrics. AUROC is widely used for assessing zero-shot detection methods [27] because it considers the True Positive Rate (TPR) and False Positive Rate (FPR) across different classification thresholds. This makes AUROC particularly useful for evaluating detector performance at different thresholds. The $F_{1}\text{ Score}$ provides a comprehensive evaluation of detector capabilities by balancing the Precision and Recall. Additionally, we provide detailed Precision and Recall scores in Appendix F for further reference, with a specific focus on Recall to highlight the detectors’ effectiveness in identifying LLM-generated text.
@@ 3 Experiments and discussion
In this section, we organize our experiments and discussions from five distinct perspectives: (1) Benchmarking the cutting-edge detectors: We evaluate the current SOTA detectors against our benchmark to identify ongoing challenges. (2) Robustness analysis: We analyze the factors contributing to robustness issues across various domains, LLMs, and attack scenarios. (3) Assessing generalization: We investigate how well detectors perform on data distribution they were not specifically trained on, highlighting their out-of-distribution robustness. (4) Length discrimination: We examine the detectors’ ability to differentiate between texts of varying lengths and discuss the impact of training on such texts. (5) Real-world human writing scenarios: We analyze the effects of real-world post-processing and mistake in human-written texts, discussing their implications to provide more nuanced and valuable insights.
@@ 3.1 Benchmarking detectors
We employed a variety of SOTA detectors to assess the difficulty of DetectRL. Given that LLMs in real-world scenarios are often black-box and inaccessible, we exclude watermarking methods from our evaluation. Our evaluation encompasses prominent zero-shot techniques and supervised fine-tuned classifiers, including Log-Likelihood [28], Entropy [29], Rank [30], Log-Rank [30], LRR [31], NPR [31], DetectGPT [27], Fast-DetectGPT [32], Revise-Detect. [33], DNA-GPT [34], Binoculars [35], RoBERTa Classifier (RoB [36]), and XLM-RoBERTa Classifier [^4] Please refer to Appendix E.2 for comprehensive detectors specifications. (X-RoB [37]). For the white-box zero-shot detection method, we employ the GPT-Neo-2.7B [38] as the scoring model, in line with the methodology proposed in Fast-DetectGPT [32], to detect the text generated by black-box LLMs. For the black-box zero-shot detection method like Revise-Detect. [33] and DNA-GPT [34], we use GPT-4o-Mini [39] to perform operations such as text revision and text continuation. For supervised detectors, all classifiers are trained using the same parameters. For detailed training parameters settings, please refer to Table 12 and AppendixE.2.
@@
Table 3: The overall leaderboard for LLM-generated text detectors in real-world scenarios ranks detectors based on their robustness and generalization across various domains, LLMs, and attack scenarios. It also considers the impact of text length in training-time and test-time phase, as well as performance against real-world human writing factors.
@@
Detectors $\downarrow$ AUROC $F_{1}$ AUROC $F_{1}$ AUROC $F_{1}$ $F_{1}$ $F_{1}$ $F_{1}$ $F_{1}$ $F_{1}$ AUROC $F_{1}$
@@
Log-Likelihood - 65.35 57.55 45.68 41.32 68.00 59.38 75.84 67.22 63.71 56.36
Entropy - 48.39 29.71 67.84 57.23 39.06 20.55 28.82 02.14 46.02 27.40
Rank - 57.17 54.62 36.87 22.47 56.26 50.90 55.08 51.90 51.34 44.97
Log-Rank - 67.01 60.09 46.74 42.60 67.58 57.57 76.40 69.88 64.43 57.53
LRR - 70.54 61.34 50.09 38.38 64.65 53.09 76.61 68.99 65.47 55.45
NPR - 53.85 49.65 34.59 18.31 54.96 52.30 50.09 45.39 48.38 41.41
DetectGPT - 22.15 00.00 12.21 00.00 58.95 50.83 44.43 35.25 34.43 21.50
Revise-Detect. - 70.40 37.51 50.34 46.07 73.24 64.29 75.01 68.71 67.24 54.14
@@
Log-Likelihood - 62.89 57.80 43.32 28.10 70.03 60.73 75.65 65.90 62.97 53.13
Entropy - 46.84 23.29 52.25 30.42 45.34 16.56 43.48 66.75 46.97 34.25
Rank - 52.19 49.32 41.68 22.78 50.40 41.74 57.05 54.40 50.33 42.06
Log-Rank - 62.84 56.87 43.32 30.12 70.89 63.09 77.97 66.66 63.75 54.18
LRR - 61.61 52.12 43.30 18.91 71.17 65.51 83.65 75.51 64.93 53.01
NPR - 50.29 43.81 41.64 32.91 44.64 34.77 52.53 48.68 47.27 40.04
DetectGPT - 43.46 26.27 32.86 12.56 26.72 00.00 36.71 20.40 34.93 14.80
DNA-GPT - 61.87 55.04 48.88 25.67 71.48 60.77 75.22 62.89 64.36 51.09
Revise-Detect. - 70.10 62.72 49.87 27.28 69.84 59.03 75.65 65.87 66.36 53.72
Binoculars - 88.14 82.50 55.15 39.35 93.30 88.20 96.64 92.30 83.30 75.58
Fast-DetectGPT - 65.56 59.55 30.01 00.00 65.99 57.58 76.79 69.08 59.58 46.55
@@
Significant Challenges
Our benchmarks reveal significant challenges in the current LLM-generated text detection research. We found that incorporating a mix distribution of domains, LLMs, and attack types increases the testing pressure of zero-shot methods. For example, in the multi-LLM setting, the average AUROC of all zero-shot detectors is only 58.61%. This is because data from each LLM spans various domains and attack methods, leading to substantial distribution differences even within data from the same LLM. These variations are often overlooked in ideal testing environments, making it difficult for zero-shot detectors developed based on them to work effectively. Specifically, zero-shot detectors struggle against powerful LLMs, achieving an average AUROC of only 77.67% on texts generated via direct prompting, with only Binoculars surpassing a 90% AUROC. The performance of these detectors declines markedly under well-designed attacks that simulate real-world scenarios, with average decreases of 1.97% in prompt attacks, 15.67% in paraphrase attacks, 38.43% in perturbation attacks, and 18.17% in data mixing scenarios. In contrast, supervised methods demonstrate impressive effectiveness, achieving an average AUROC of 99.40% on data generated through direct prompting and maintaining robustness against well-designed attacks.
Unexpectedly, recent advancements in LLM-generated text detection, such as DetectGPT [27], NPR [31], Fast-DetectGPT [32], and DNA-GPT [34], did not perform as expected on our benchmark. Their performance was even weaker than some traditional zero-shot baselines. Analysis across various domains and LLMs revealed a general lack of robustness as a potential underlying issue. For instance, DetectGPT’s performance was notably low, with only 22.15% AUROC in academic writing (ArXiv) and 12.21% AUROC in news writing (Xsum), though it achieved 58.95% AUROC in creative writing and 44.43% AUROC in social media (Yelp Review). A similar trend was observed with the best zero-shot detector, Binoculars, which performed more than 10% lower in academic writing and news writing compared to other domains. Additionally, Binoculars showed significantly reduced effectiveness on text generated by Claude, achieving only 55.15% AUROC, while presenting 88.14%, 93.30%, and 96.64% AUROC on text generated by GPT-3.5, PaLM-2, and Llama-2, respectively. These findings suggest that the performance differences of detectors across different domains and LLMs become significantly more pronounced when subjected to well-designed attacks.